# Deep Reinforcement Learning for Wireless Scheduling in Distributed Networked Control

### v4

@@ Abstract
G. Pang and K. Huang contributed equally to the work. G. Pang, B. Vucetic, Y. Li, and W. Liu are with School of Electrical and Computer Engineering, The University of Sydney, Australia. Emails: {gaoyang.pang; branka.vucetic; yonghui.li; wanchun.liu}@sydney.edu.au. K. Huang is with Huawei Shanghai Research Center, Shanghai, China. Email: huangkang9@huawei.com. D. E. Quevedo is with the School of Electrical Engineering and Robotics, Queensland University of Technology (QUT), Brisbane, Australia. Email: dquevedo@ieee.org.
@@ I Introduction
The Fourth Industrial Revolution, Industry 4.0, is the automation of conventional manufacturing and industrial processes through flexible mass production. Automatic control in Industry 4.0 requires large-scale and interconnected deployment of massive spatially distributed industrial devices, such as sensors, actuators, machines, robots, and controllers. Eliminating the communication wires is a game-changer in reshaping traditional factories. In this regard, wireless networked control has become one of the most important technologies of Industry 4.0. It offers high-scalable and low-cost deployment capabilities and enables many industrial applications, such as smart city, smart manufacturing, smart grids, e-commerce warehouses and industrial automation systems [1, 2].
@@
Unlike traditional cable-based networked control systems, in a large-scale wireless networked control system (WNCS), communication resources are limited. Not surprisingly, during the past decade, state estimation and transmission scheduling in WNCS have drawn a lot of attention in the research community. For example, some recent works have investigated the estimation problem under dynamic sampling periods [3] and random access protocol scheduling [4]. Focusing on the sensor-controller communications only (the controller-actuator co-located scenario), the optimal transmission scheduling problem of over a single frequency channel for achieving the best remote estimation quality was extensively investigated in [5, 6, 7, 8]. For the multi-frequency channel scenario, the optimal scheduling policy and structural results were obtained in [9]. The optimal transmission power scheduling problem of an energy-constrained remote estimation system was studied in [10]. The joint scheduling and power allocation problems of multi-plant-multi-frequency WNCSs were investigated in [11, 12] for achieving the minimum overall transmission power consumption. From a cyber-physical security perspective, denial-of-service (DoS) attackers’ scheduling problems were investigated in [13, 14], for maximally deteriorating the WNCSs’ performance.
@@
In the WNCS examined in the above works, the scheduling problem is commonly reformulated into a Markov decision process (MDP) problem. The MDP problem can be solved using traditional algorithms (e.g., value iteration [9, 2, 15], policy iteration [16], and linear programming [17, 18]) and reinforcement learning algorithms (e.g., Q-learning [19] and deep Q-Networks [20, 21]). Traditional algorithms and Q-learning are effective when the system scale is small. For example, numerical results of the optimal scheduling of a two-sensor-one-frequency using value iteration system were presented in [9]. These methods involve operations over every possible state and action. As the number of actions increases, the number of state-action pairs grows, leading to a dramatic increase in the computations required. For example, when the action space is large, Q-learning necessitates constructing a Q-value table, which demands substantial storage proportional to the size of the action space, leading to the curse of dimensionality [19, 22, 23]. Deep Q-Networks (DQN) address this issue using deep neural networks (DNNs) for value function approximations, thus eliminating the need for extensive Q-value tables. Some recent works [20, 21] have applied DQN to solve multi-system-multi-frequency scheduling problems in different WNCS scenarios.
@@
However, handling large action spaces also presents significant challenges in DQN. The size of the DNNs required for DQN expands dramatically with respect to the action space. This can hinder effective exploration of the action space during training, resulting in increased storage and computational requirements. Ma et al. [24] proposed decoupling the original problem into subproblems and developing multiple deep reinforcement learning (DRL) algorithms to solve the corresponding subproblems with reduced decision space. However, the presence of multiple learning agents creates a non-stationary environment, which can destabilize learning and make convergence more difficult. To deal with large decision spaces, recent works resort to stochastic policy-based DRL algorithms [25, 26, 27]. A stochastic policy refers to a policy that specifies a probability distribution over actions, given the current state of the environment. Ni et al. [25] leveraged a Graph Neural Network (GNN) based DRL to learn stochastic policies, i.e., Proximal Policy Optimization (PPO) for warehouse scheduling. Yang et al. [26] developed an actor-critic DRL for learning stochastic policies with continuous action space for scheduling, power allocation, and modulation scheme adaptation. However, it has been proved that the optimal policy is deterministic for unconstraint MDP problems [28]. Here, a deterministic policy means that the action in a given state is determined by a fixed function without any randomness. Consequently, deterministic policy-based DRL methods are generally more preferable.
@@
We note that most of the existing works only focus on WNCSs which are only partially distributed [5, 6, 7, 9, 11, 12, 20, 21]. For example, Chen et al. [29] focus on the sensor scheduling in the uplink transmission of a WNCS, using either DQN or Deep Deterministic Policy Gradient (DDPG). Redder et al. [30] developed a DQN-based algorithm for optimal actuator scheduling in the downlink transmission of a WNCS over a Markov fading channel. Existing works only consider optimizing either the uplink or downlink transmission for WNCSs. In a fully distributed setting, both downlink (controller-actuator) and uplink (sensor-controller) transmissions are crucial for stabilizing each plant. The joint uplink and downlink scheduling problem of fully distributed multi-plant WNCS has never been considered in the open literature.
@@
  • We propose a distributed $N$-plant-$M$-frequency WNCS model, where the controller schedules the uplink and downlink transmissions of all $N$ plants and the spatial diversity of different communication links are taken into account. The controller generates sequential predictive control commands for each of the plants based on pre-designed deadbead control laws. [^2] Note that deadbead controller is commonly considered as a time-optimal controller that takes the minimum time for setting the current plant state to the origin. Different from uplink or downlink only scheduling, a joint scheduling algorithm has a larger action space and also needs to automatically balance the tradeoff between the uplink and the downlink due to the communication resource limit. To the best of our knowledge, joint uplink and downlink transmission scheduling of distributed WNCSs has not been investigated in the open literature.
@@
  • We construct a finite-length countable vector state of the WNCS in terms of the time duration between consecutive received packets at the controller and the actuators. Then, we prove that the per-step cost function of the WNCS is determined by the time-duration-related vector state. Building on this, we formulate the optimal transmission scheduling problem into a Markov decision process (MDP) problem with an countable state space for achieving the minimum expected total discounted cost. We propose effective action space reduction and action embedding methods that can be applied to DRL algorithms, including foundational ones like DQN and more advanced ones such as DDPG and Twin Delayed Deep Deterministic Policy Gradient (TD3), for solving the problem. Numerical results illustrate that the proposed algorithm can reduce the expected cost significantly compared to available benchmark policies.
@@ II Distributed WNCS with Shared Wireless Resource
We consider a distributed WNCS system with $N$ plants and a central controller as illustrated in Fig. 1. The output of plant $i$ is measured by smart sensor $i$, which sends pre-filtered measurements (local state estimates) to the controller. The controller applies a remote (state) estimator and a control algorithm for plant $i$. It then generates and sends a control signal to actuator $i$, thereby closing the loop. A key limitation of the WNCS is that smart sensor $i$ cannot communicate with actuator $i$ directly due to the spatial deployment. Instead, the uplink (sensor-controller) and downlink (controller-actuator) communications for the $N$ plants share a common wireless network with only $M$ frequency channels, where $M<2N$. Thus, not every node is allowed to transmit at the same time and communications need to be scheduled. As shown in Fig. 1, we shall focus on a setup where scheduling is done at the controller side, which schedules both the downlink and uplink transmissions. [^3] Such a setup is practical, noting that most of the existing wireless communication systems including 5G support both uplink and downlink at base stations [31].
Each smart sensor has three modules: remote estimator, control algorithm and local Kalman filter. The first two modules are copies of the ones at the controller to reconstruct the control input of the plant. Such information is sent to the Kalman filter for accurate local state estimation. The details will be presented in the sequel.
@@ II-B Smart Sensors
Due to the noise measurement, each smart sensor runs a Kalman filter to estimate the current plant state as below [33]: [^4] In this subsection, we focus on smart sensor $i$ and the index $i$ of each quantity is omitted for clarity
@@
where $\mathbf{x}^{s}_{k|k-1}$ and $\mathbf{x}^{s}_{k}$ are the prior and posterior state estimation at time $k$, respectively, and $\mathbf{P}^{s}_{k|k-1}$ and $\mathbf{P}^{s}_{k}$ are the (estimation error) covariances of $\mathbf{e}^{s}_{k|k-1}\triangleq\mathbf{x}_{k}-\mathbf{x}^{s}_{k|k-1}$ and $\mathbf{e}^{s}_{k}\triangleq\mathbf{x}_{k}-\mathbf{x}^{s}_{k}$, respectively. $\mathbf{K}_{k}$ is the Kalman gain at time $k$. We assume that each $(\mathbf{A},\mathbf{C})$ is observable and $(\mathbf{A},\mathbf{Q}^{w})$ is controllable [9]. Thus, the Kalman gain $\mathbf{K}_{k}$ and error covariance matrix $\mathbf{P}^{s}_{k}$ converge to constant matrices $\hat{\mathbf{K}}$ and $\hat{\mathbf{P}}^{s}$, respectively, i.e., the smart sensor is in the stationary mode. As foreshadowed, the smart sensor employs a remote estimator and knowledge of the control algorithm that is applied at the controller side to obtain the control input in (4). Thus, the Kalman filter (4) is the optimal estimator of the linear system (1) in terms of the estimation mean-square error [34]. The remote estimator and control algorithm will be presented in Sections II-C and II-D, respectively.
As shown in Fig. 1, at every scheduling instant $k$, the smart sensor sends the local estimate ${\mathbf{x}}_{k}^{s}$ (rather than the raw measurement $\mathbf{y}_{k}$) to the controller.
@@ II-C Remote Estimation
From (9), the controller’s current estimation depends on the most recently received sensor estimation. Let $\tau_{i,k}\in\{1,2,\cdots\}$ denote the age-of-information (AoI) (see [15] and reference therein) of sensor $i$’s packet observed at time $k$, i.e., the number of elapsed time slots from the latest successfully delivered sensor $i$’s packet before the current time $k$, which reflects how old the most recently received sensor measurement is. Then, it is easy to see that the updating rule of $\tau_{i,k}$ is given by
@@ II-D Control Algorithm
Due to the fact that downlink transmissions are unreliable, actuator $i$ may not receive the controller’s control-command-carrying packets, even when transmissions are scheduled. We adopt a predictive control approach [32, 35] to provide robustness against packet failures: the controller sends a length-$v_{i}$ sequence of control commands including both the current command and the predicted future commands to the actuator once scheduled; if the current packet is lost, the actuator will apply the previously received predictive command for the current time slot as the control input.
@@
It can be verified that if the current state estimation $\hat{\mathbf{x}}_{i,k}$ is perfect and the plant $i$ is disturbance free, then the plant state $\mathbf{x}_{i,k}$ would be set to zero after applying all $v_{i}$ steps of the control sequence in (13). Such a deadbeat controller is commonly considered as a time-optimal controller that takes the minimum time for setting the current plant state to the origin [36]. We note that the deadbeat control law may not be cost-optimal optimal to minimize the quadratic cost function defined in Section IV, and the optimal control law may depend on the scheduling policy. Since the current work focuses on the transmission scheduling design of the $N$-plant-$M$-frequency WNCS, the optimal joint control-scheduling problem can be investigated in our future work. Specifically, we aim to identify an optimal scheduling policy that minimizes the quadratic cost for a given deadbeat controller.
@@
Let $\eta_{i,k}\in\{1,2,\cdots\}$ denote the AoI of controller’s packet at actuator $i$ observed at time $k$, i.e., the number of passed time slots (including the current time slot) since the latest received control packet by actuator $i$. Its updating rule is given as
@@ II-E Communication Scheduler
The packet transmissions of each link are modeled as i.i.d. packet dropout processes. Unlike most of the existing works wherein transmission scheduling of WNCSs assumes that different node transmissions at the same frequency channel have the same packet drop probability [20], we here consider a more practical scenario by taking into account the spatial diversity of different transmission nodes – each frequency has different dropout probabilities for different uplink and downlink transmissions. The packet success probabilities of the uplink and downlink of plant $i$ at frequency $m$ are given by $\xi^{s}_{m,i}$ and $\xi^{c}_{m,i}$, respectively, where
@@
Acknowledgment feedback. We note that the Transmission Control Protocol (TCP) is commonly adopted in commercial telecommunication systems for data transmission, such as 5G and WiFi [39, 40, 41]. TCP relies on acknowledgment feedback, where the receiver sends a one-bit acknowledgment signal back to the sender to confirm the successful receipt of data packets. This feedback mechanism ensures reliable data transmission by verifying the delivery of data. Therefore, we assume that both the uplink and downlink transmissions adopt the acknowledgment feedback scheme. In particular, the actuator sends one-bit feedback signal of $\gamma_{i,k}$ to the controller, and the controller sends $\beta_{i,k}$ as well as the received $\gamma_{i,k}$ to sensor $i$ each time with negligible overhead. From $\{\beta_{i,k}\}$ and (9), the smart sensor knows the estimated plant state $\hat{\mathbf{x}}_{i,k}$ by the controller; then, using $\hat{\mathbf{x}}_{i,k}$, $\{\gamma_{i,k}\}$ and (18), the sensor can calculate the applied control input $\mathbf{u}_{i,k}$, which is utilized for local state estimation as mentioned in (4). Note that sending one-bit acknowledgement feedback leads to negligible overhead in contrast to sending the applied control input $\mathbf{u}_{i,k}$ from the controller to the sensors.

### v5

@@ I Introduction
The Fourth Industrial Revolution, Industry 4.0, is the automation of conventional manufacturing and industrial processes through flexible mass production. Automatic control in Industry 4.0 requires large-scale and interconnected deployment of massive spatially distributed industrial devices, such as sensors, actuators, machines, robots, and controllers. Eliminating the communication wires is a game-changer in reshaping traditional factories. In this regard, wireless networked control has become one of the most important technologies of Industry 4.0. It offers highly scalable and low-cost deployment capabilities and enables many industrial applications, such as smart cities, smart manufacturing, smart grids, e-commerce warehouses and industrial automation systems [1, 2].
@@
Unlike traditional cable-based networked control systems, in a large-scale wireless networked control system (WNCS), communication resources are limited. Not surprisingly, during the past decade, state estimation and transmission scheduling in WNCS have drawn a lot of attention in the research community. For example, some recent works have investigated the estimation problem under dynamic sampling periods [3] and random access protocol scheduling [4]. Focusing on the sensor-controller communications only (the controller-actuator co-located scenario), the optimal transmission scheduling problem over a single frequency channel for achieving the best remote estimation quality was extensively investigated in [5, 6, 7, 8]. For the multi-frequency channel scenario, the optimal scheduling policy and structural results were obtained in [9]. The optimal transmission power scheduling problem of an energy-constrained remote estimation system was studied in [10]. The joint scheduling and power allocation problems of multi-plant-multi-frequency WNCSs were investigated in [11, 12] for achieving the minimum overall transmission power consumption. From a cyber-physical security perspective, the scheduling and control problems posed by denial‑of‑service (DoS) attackers were investigated in [13, 14, 15], for maximally deteriorating the WNCSs’ performance.
@@
In the WNCS examined in the above works, the scheduling problem is commonly reformulated into a Markov decision process (MDP) problem. The MDP problem can be solved using traditional algorithms (e.g., value iteration [9, 2, 16], policy iteration [17], and linear programming [18, 19]) and reinforcement learning algorithms (e.g., Q-learning [20, 21] and deep Q-Networks [22, 23]). Traditional algorithms and Q-learning are effective when the system scale is small. For example, numerical results of the optimal scheduling of a two-sensor-one-frequency system using value iteration were presented in [9]. These methods involve operations over every possible state and action. As the number of actions increases, the number of state-action pairs grows, leading to a dramatic increase in the computations required. For example, when the action space is large, Q-learning necessitates constructing a Q-value table, which demands substantial storage proportional to the size of the action space, leading to the curse of dimensionality [20, 24, 25]. Deep Q-Networks (DQN) address this issue using deep neural networks (DNNs) for value function approximations, thus eliminating the need for extensive Q-value tables. Some recent works [22, 23, 26] have applied DQN to solve multi-system-multi-frequency scheduling problems in different WNCS scenarios.
@@
However, handling large action spaces also presents significant challenges in DQN. The size of the DNNs required for DQN expands dramatically with respect to the action space. This can hinder effective exploration of the action space during training, resulting in increased storage and computational requirements. Ma et al. [27] proposed decoupling the original problem into subproblems and developing multiple deep reinforcement learning (DRL) algorithms to solve the corresponding subproblems with reduced decision space. However, the presence of multiple learning agents creates a non-stationary environment, which can destabilize learning and make convergence more difficult. To deal with large decision spaces, recent works resort to stochastic policy-based DRL algorithms [28, 29, 30]. A stochastic policy refers to a policy that specifies a probability distribution over actions, given the current state of the environment. Ni et al. [28] leveraged a Graph Neural Network (GNN) based DRL to learn stochastic policies, i.e., Proximal Policy Optimization (PPO) for warehouse scheduling. Yang et al. [29] developed an actor-critic DRL for learning stochastic policies with a continuous action space for scheduling, power allocation, and modulation scheme adaptation. However, it has been proved that the optimal policy is deterministic for unconstrained MDP problems [31]. Here, a deterministic policy means that the action in a given state is determined by a fixed function without any randomness. Consequently, deterministic policy-based DRL methods are generally preferable.
@@
We note that most of the existing works focus on WNCSs which are partially distributed [5, 6, 7, 9, 11, 12, 22, 23]. For example, Chen et al. [32] focus on the sensor scheduling in the uplink transmission of a WNCS, using either DQN or Deep Deterministic Policy Gradient (DDPG). Redder et al. [33] developed a DQN-based algorithm for optimal actuator scheduling in the downlink transmission of a WNCS over a Markov fading channel. Wang et al. [34] also developed a GNN-based multi-agent DRL for optimal actuator scheduling in a WNCS. Existing works consider only optimizing either the uplink or downlink transmission for WNCSs. In a fully distributed setting, both downlink (controller-actuator) and uplink (sensor-controller) transmissions are crucial for stabilizing each plant. The joint uplink and downlink scheduling problem of fully distributed multi-plant WNCS has never been considered in the open literature.
@@
  • We propose a distributed $N$-plant-$M$-frequency WNCS model, where the controller schedules the uplink and downlink transmissions of all $N$ plants and the spatial diversity of different communication links is taken into account. The controller generates sequential predictive control commands for each of the plants based on pre-designed deadbeat control laws. [^1] Note that deadbeat controller is commonly considered as a time-optimal controller that takes the minimum time for setting the current plant state to the origin. Different from uplink or downlink only scheduling, a joint scheduling algorithm has a larger action space and also needs to automatically balance the trade-off between the uplink and the downlink due to the communication resource limit. To the best of our knowledge, joint uplink and downlink transmission scheduling of distributed WNCSs has not been investigated in the open literature.
@@
  • We construct a finite-length countable vector state of the WNCS in terms of the time duration between consecutive received packets at the controller and the actuators. Then, we prove that the per-step cost function of the WNCS is determined by the time-duration-related vector state. Building on this, we formulate the optimal transmission scheduling problem into an MDP problem with a countable state space for achieving the minimum expected total discounted cost. We propose effective action space reduction and action embedding methods that can be applied to DRL algorithms, including foundational ones like DQN and more advanced ones such as DDPG and Twin Delayed Deep Deterministic Policy Gradient (TD3), for solving the problem. Numerical results illustrate that the proposed algorithm can reduce the expected cost significantly compared to available benchmark policies.
@@ II Distributed WNCS with Shared Wireless Resource
We consider a distributed WNCS system with $N$ plants and a central controller as illustrated in Fig. 1. Each smart sensor employs a local Kalman filter. The output of plant $i$ is measured by smart sensor $i$, which sends pre-filtered measurements (local state estimates) to the controller. The central controller applies a remote (state) estimator and a control algorithm for plant $i$. It then generates and sends a control signal to the actuator $i$, thereby closing the loop. The uplink (sensor-controller) and downlink (controller-actuator) communications for the $N$ plants share a common wireless network with only $M$ frequency channels, where $M<2N$. Thus, not every node is allowed to transmit at the same time and communications need to be scheduled. As shown in Fig. 1, we shall focus on a setup where scheduling is done at the controller side, which schedules both the downlink and uplink transmissions. [^2] Such a setup is practical, noting that most of the existing wireless communication systems, including 5G, support both uplink and downlink at base stations [35].
@@ II-B Smart Sensors
Due to the measurement noise, each smart sensor runs a Kalman filter to estimate the current plant state as below [37]: [^3] In this subsection, we focus on smart sensor $i$ and the index $i$ of each quantity is omitted for clarity.
@@
where $\mathbf{x}^{s}_{k|k-1}$ and $\mathbf{x}^{s}_{k}$ are the prior and posterior state estimation at time $k$, respectively, and $\mathbf{P}^{s}_{k|k-1}$ and $\mathbf{P}^{s}_{k}$ are the (estimation error) covariances of $\mathbf{e}^{s}_{k|k-1}\triangleq\mathbf{x}_{k}-\mathbf{x}^{s}_{k|k-1}$ and $\mathbf{e}^{s}_{k}\triangleq\mathbf{x}_{k}-\mathbf{x}^{s}_{k}$, respectively. $\mathbf{K}_{k}$ is the Kalman gain at time $k$. We assume that each $(\mathbf{A},\mathbf{C})$ is observable and $(\mathbf{A},\mathbf{Q}^{w})$ is controllable [9]. Thus, the Kalman gain $\mathbf{K}_{k}$ and error covariance matrix $\mathbf{P}^{s}_{k}$ converge to constant matrices $\hat{\mathbf{K}}$ and $\hat{\mathbf{P}}^{s}$, respectively, i.e., the smart sensor is in the stationary mode. As shown in Fig. 1, the smart sensor is co-located with the actuator in the plant [38] and can obtain the control input in (4). Thus, the Kalman filter (4) is the optimal estimator of the linear system (1) in terms of the estimation mean-square error [39].
As shown in Fig. 1, at every scheduling instant $k$, the smart sensor sends the local estimate ${\mathbf{x}}_{k}^{s}$ (rather than the raw measurement $\mathbf{y}_{k}$) to the controller [37].
@@ II-C Remote Estimation
From (9), the controller’s current estimation depends on the most recently received sensor estimation. Let $\tau_{i,k}\in\{1,2,\cdots\}$ denote the age-of-information (AoI) (see [16] and reference therein) of sensor $i$’s packet observed at time $k$, i.e., the number of elapsed time slots since the latest successfully delivered sensor $i$’s packet before the current time $k$, which reflects how old the most recently received sensor measurement is. Then, it is easy to see that the updating rule of $\tau_{i,k}$ is given by
@@ II-D Control Algorithm
Due to the fact that downlink transmissions are unreliable, actuator $i$ may not receive the controller’s control-command-carrying packets, even when transmissions are scheduled. We adopt a predictive control approach [36, 40] to provide robustness against packet failures: the controller sends a length-$v_{i}$ sequence of control commands including both the current command and the predicted future commands to the actuator once scheduled; if the current packet is lost, the actuator applies the previously received predictive command as the control input for the current time slot.
@@
It can be verified that if the current state estimation $\hat{\mathbf{x}}_{i,k}$ is perfect and the plant $i$ is disturbance free, then the plant state $\mathbf{x}_{i,k}$ would be set to zero after applying all $v_{i}$ steps of the control sequence in (13). Such a deadbeat controller is commonly considered as a time-optimal controller that takes the minimum time for setting the current plant state to the origin [41]. We note that the deadbeat control law may not be cost-optimal to minimize the quadratic cost function defined in Section IV, and the optimal control law may depend on the scheduling policy. Since the current work focuses on the transmission scheduling design of the $N$-plant-$M$-frequency WNCS, the optimal joint control-scheduling problem can be investigated in our future work. Specifically, we aim to identify an optimal scheduling policy that minimizes the quadratic cost for a given deadbeat controller.
@@
Let $\eta_{i,k}\in\{1,2,\cdots\}$ denote the AoI of the controller’s packet at actuator $i$ observed at time $k$, i.e., the number of elapsed time slots (including the current time slot) since the actuator $i$ last received a control packet. Its updating rule is given as
@@ II-E Communication Scheduler
The packet transmissions of each link are modeled as i.i.d. packet dropout processes. Unlike most of the existing works wherein transmission scheduling of WNCSs assumes that transmissions from different nodes on the same frequency channel have the same packet drop probability [22], we here consider a more practical scenario by taking into account the spatial diversity of different transmission nodes – each frequency has different dropout probabilities for different uplink and downlink transmissions. The packet success probabilities of the uplink and downlink of plant $i$ at frequency $m$ are given by $\xi^{s}_{m,i}$ and $\xi^{c}_{m,i}$, respectively, where
@@
Acknowledgment feedback. We note that the Transmission Control Protocol (TCP) is commonly adopted in commercial telecommunication systems for data transmission, such as 5G and WiFi [44, 45, 46]. TCP relies on acknowledgment feedback, where the receiver sends a one-bit acknowledgment signal back to the sender to confirm the successful receipt of data packets. This feedback mechanism ensures reliable data transmission by verifying the delivery of data. Therefore, we assume that both the uplink and downlink transmissions adopt the acknowledgment feedback scheme. In particular, the actuator sends a one-bit feedback signal of $\gamma_{i,k}$ to the controller, and the controller sends $\beta_{i,k}$ as well as the received $\gamma_{i,k}$ to sensor $i$ each time with negligible overhead. From $\{\beta_{i,k}\}$ and (9), the smart sensor knows the estimated plant state $\hat{\mathbf{x}}_{i,k}$ by the controller; then, using $\hat{\mathbf{x}}_{i,k}$, $\{\gamma_{i,k}\}$ and (18), the sensor can calculate the applied control input $\mathbf{u}_{i,k}$, which is utilized for local state estimation as mentioned in (4). Note that sending one-bit acknowledgment feedback leads to negligible overhead in contrast to sending the applied control input $\mathbf{u}_{i,k}$ from the controller to the sensors.